If y varies inversely as x with a exponent 2 and y=2 when x=4, what is the constant of variation?
Answer by stanbon(75887)

If y varies inversely as x with a exponent 2 and y=2 when x=4, what is the constant of variation?
--------------------------------
y = k/x^2
---
Solve for "k" using y=2 when x = 4
---
2 = k/4^2
---
k = 2*16
k = 32
===========================
Cheers,
Stan H.